Given that John is two years older than eric and the sum of whir ages is 26 years how old is Eric and john
Answer by ikleyn(52787)

.
Eric + John  = 26


E    + (E+2) = 26

2E = 26 - 2

2E = 24

E = 24/2 = 12;    J = E+2 = 12+2 = 14.


Answer.  12 and 14 years old.